If water has a surface tension of 7 x 10^2 Nm and an angle of contact with water of zero. At what height does it rise in a capillary of diameter 0.5 mm?
What are the dimensions of surface tension?
1
Expert's answer
2020-10-21T09:54:17-0400

As the height of rise of liquid in capillary is equal to: "h=\frac{2\times σ\times \cos \theta}{(p\times g\times R)}"

Where: "\u03c3" -surface tension coefficient, "σ = 7\times10^2" H/m, p is the fluid density, "p" (water) "= 10^3" kg/m"^3" , "g" -acceleration of free fall, "g = 10" m/S"^2" , "R" is the radius of the capillary "R =\frac{1}{2}\times 5\times10^{-4}" m.

Then: "h=\frac{2\times σ\times \cos \theta }{(p\times g\times R)}=\frac{2\times7\times10^2 \times \cos 0}{(103\times 10 \times \frac{1}{2}\times5\times10^{-4})}=5600" m

The size of the surface tension is

"S= h\times \pi \times R^2 = 5600\times" "\pi \times (2.5\times 10^{-4})^2\approx 0.001" m^2.